# A New Approach for Fourier Extension Based on Weighted Generalized Inversethanks: The research is partially supported by National Natural Science Foundation of China (Nos. 12261131494, RSF-NSFC 23-41-00002)

### v2

@@ Abstract
This paper examines the Fourier extension from a new perspective of solving the compact operator equation with perturbed data. By converting the approximation target from the best approximate solution to the weighted best approximate solution, the oscillation in the extended region has been overcome. The error estimation of the solution is theoretically established. Furthermore, we point out the difficulties faced by the original weighted operator in calculation due to the limitation of machine precision and propose an effective correction operator. The relevant parameters involved in the method are further tested, and finally the effectiveness of the method is verified through numerical experiments.
@@
Fourier Extensions, Fourier Continuation, Weighted Generalized Inverse
{AMS}
42A10, 65T40, 65T50
@@ 1 Introduction
Due to its flexibility, frame approximation has received increasing attention in recent years. Its related theories and numerical algorithms are being widely studied and have made substantial progress [2, 4, 3, 5, 12, 8]. Fourier extension is an important form of frame approximation, which has significantly contributed to the advancement of the field of frame theory. Initially proposed to mitigate the Gibbs phenomenon observed in classical Fourier series, Fourier extension has since witnessed extensive research. In recent years, many scholars have made substantial contributions to developing its mathematical foundations, devising efficient algorithms, and exploring its applications across various fields [6, 13, 14, 16, 15, 17, 7, 20].
Similar to general frame approximations, Fourier extension also encounters challenges due to its near-redundancy. Specifically, it faces the issue of ill-conditioning in the least-squares system used to compute the coefficients of the approximation. Fortunately, accurate Fourier extension approximation can still obtained via suitable regularization. Currently, Truncated Singular Value Decomposition (TSVD) is the commonly used regularization technique, and various fast algorithms have been developed on this basis [14, 16, 21]. However, the solutions calculated by existing methods exhibit oscillations in the extended region, which is detrimental to the application of the extension method in solving partial differential equations (PDEs). A post-processing method described in the literature [15]demonstrates that it is possible to find extension functions with desirable properties without compromising accuracy.
According to the extension theorem of Sobolev spaces, it is possible to identify extension functions with desirable properties. This raises a question: Can we find an effective method to directly calculate such extension functions? In fact, when $N$ is finite, $\{e^{\text{i}\frac{k\pi}{T}t}\}_{k=-N}^{N}$ is a basis on [-1,1], that is, the solution to the least squares problem is unique. However, when the TSVD is used, since the singular vectors with singular values $\sigma_{k}\leq\epsilon$ (where $\epsilon$ is the truncation parameter) are discarded, the problem actually becomes a multi-solution. TSVD only yields the solution which minimizes the Euclidean norm of the Fourier coefficients. When a function has a certain smoothness, its Fourier coefficient has a corresponding decay. The smoother the function, the faster the decay rate. Therefore, it is natural to restrict its decay in the selection of solutions. The idea of enforcing special features of regularization approximations has a long history. In his original paper [19], Tikhonov (1963) defined the regularization solution $x_{\alpha}$ of the equation
@@
where $\mathcal{L}$ is a differential operator. Along similar lines, Hansen et. al developed GTSVD and MTSVD methods [9, 11, 10] and Natterer (1984) introduced “regularization in Hilbert scales” [18]. Interestingly, Natterer (1984) has shown that “oversmoothing” (the unknown solution fails to have the expected smoothness) makes no trouble, that is, we can impose stronger constraints than the smoothness of the solution.
@@ 2.1 A new perspective on Fourier extension
From [13], this system is a tight frame with frame bound $T$. For $f\in H^{s}(\Lambda_{1})$, we consider the compact operator equation
@@
  • From the classical regularization theory, we know that the regularization parameter should be related to the error level $\|f-f_{N}\|_{L^{2}(\Lambda_{1})}$, which explains “the coefficient norm first increases exponentially fast, before decaying to $\mathcal{O}(1)$ in the limit.” mentioned in [5]. When $N$ is small, the error $\|f-f_{N}\|_{L^{2}(\Lambda_{1})}$ is large, and $\epsilon$ is too small for the regularization parameter required for stable solution. As $N$ increases, the error gradually decreases, and after a certain threshold, $\epsilon$ can meet the requirements of stable solution.Therefore, we can actually draw a conclusion: the solutions with large coefficients are unimportant because they mean that $\|f-f_{N}\|_{L^{2}(\Lambda_{1})}$ is large, that is, a desirable approximation to $f$ has not been obtained.
@@
    $\|f-\mathcal{Q}_{N,1}^{\epsilon}f\|\leq\|f-f_{N}\|_{L^{2}(\Lambda_{1})}+\epsilon\|{s}_{N}\|$
@@
    $\|f-\mathcal{Q}_{N,2}^{\epsilon}f\|\leq\|f-f_{N}\|_{L^{2}(\Lambda_{1})}+\sqrt{\epsilon}\|{s}_{N}\|.$
    That is to say, if we solve the normal equation, we can only obtain an accuracy of $O(\sqrt{\epsilon})$ instead of $O({\epsilon})$. The literature [2] also shows that this is true for general frame approximations. Therefore, we should directly focus on the discretization of equation (2) rather than its normal equation.
@@ 2.2 Fourier extension towards weighted best-approximation solution
Because $f\in H^{s}(\Lambda_{1})$, we can expect the coefficient $\{\hat{c}_{k}\}$ has a corresponding decay property. Therefore, the following weighted best-approximation solution ${c}^{{\dagger}}_{\mathcal{L}}$ is a more reasonable goal:
@@
$(\mathcal{L}{c})_{k}=\sqrt{1+|k|^{2p}}\hat{c}_{k}.$ (9)
with $p\geq 0$.
@@
  • If $p=0$, then we have ${c}^{{\dagger}}={c}^{{\dagger}}_{\mathcal{L}}.$
  • If possible, we would like to take $p=s$, that is, we hope that the solution to the problem has the same smoothness as the function $f$. However, in practice, the value of $s$ may be unknown. Furthermore, if the value of $p$ is too large, the calculation will face the issue of overflow. Therefore, even if $f$ is analytic, we only take a moderate value of $p$ in the calculation. In fact, the above operator $\mathcal{L}$ faces the difficulty of numerical realization due to the limitation of machine precision. We will discuss this in section 3.2 and propose a modified operator.
  • If $p>s$, that is , the smoothness of the solution is lower than the requirement of the operator $\mathcal{L}$, i. e., no solution of (1) is an element of $\mathcal{D}(\mathcal{L})$, this is called oversmoothing. The results in the literature [18] have shown that the regularization method can still obtain the optimal convergence order in the case of oversmoothing (when the regularization parameters are reasonably selected).
Now we set the approximation target the weighted generalized solution ${c}^{{\dagger}}_{\mathcal{L}}$. At this time, the approximation of the solution can be achieved by the “regularization in the Hilbert scale”[18]. In fact, if we let
$\psi_{k}(x)=\frac{1}{\sqrt{1+k^{2p}}}\phi_{k}(x),\quad\hbox{and}\quad\Psi=\{\psi_{k}\}_{k\in\mathbb{Z}},$
@@
${c}^{{\dagger}}_{\mathcal{L}}=\mathcal{L}^{-1}\tilde{c}^{{\dagger}}.$
@@
$\mathcal{L}_{N}:\mathbb{C}^{2N+1}\rightarrow\mathbb{C}^{2N+1},{c}_{N}=\{\hat{c}_{k}\}_{|k|\leq N}\mapsto\tilde{c}_{N}=\{\sqrt{1+k^{2p}}\hat{c}_{k}\}_{|k|\leq N},$
@@
~$\tilde{\mathcal{F}}_{N}=\mathcal{F}_{N}\mathcal{L}^{-1}_{N},\quad{s}_{\mathcal{L},N}=\mathcal{P}_{N}{c}_{\mathcal{L}}^{{\dagger}}\quad\tilde{s}_{\mathcal{L},N}=\mathcal{L}_{N}{s}_{\mathcal{L},N},\quad f_{\mathcal{L},N}=\tilde{\mathcal{F}}_{N}\tilde{s}_{\mathcal{L},N}.$
@@
$c_{\mathcal{L},N}^{\epsilon}=\mathcal{L}_{N}^{-1}\tilde{c}_{N}^{\epsilon}$ (13)
@@
$\|f-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f\|_{L^{2}(\Lambda_{1})}\leq\|f-f_{\mathcal{L,N}}\|_{L^{2}(\Lambda_{1})}+\epsilon\|\tilde{s}_{\mathcal{L},N}\|.$ (15)
Proof 2.2.
@@ 3.1 Discretization based on uniform sampling
Selecting different discrete nodes will affect the discrete accuracy of the operator $\mathcal{F}_{N}$, thus having a substantial impact on the calculation results. Previous work distinguishes between uniform sampling, and an optimal sampling set called mapped symmetric Chebychev nodes [1]. In this paper, we focus on the case where the sampling points are uniform:
@@
$\left({\bf F}_{\gamma,N}^{T,p}\right)_{\ell,k}=\frac{1}{\sqrt{L}}\phi_{k}(t_{\ell})$
and ${\bf L}_{N}$ be a diagonal matrix with
$({\bf d})_{k}={\sqrt{1+|k|^{2p}}},\quad k=-N,\ldots,N.$ (18)
@@
$\tilde{\bf F}_{\gamma,N}^{T,p}={\bf F}_{\gamma,N}^{T,p}{\bf L}_{N}^{-1}$ (19)
@@
${\bf c}_{{\bf L},N}^{\epsilon}=\mathcal{L}_{N}^{-1}\tilde{\bf c}_{N}^{\epsilon},\quad{\bf Q}_{{\bf L},N}^{\epsilon}{\bf f}={\bf F}_{\gamma,N}^{T,p}{\bf c}_{{\bf L},N}^{\epsilon}.$ (22)
@@ 3.2 The influence of the parameter $p$ on calculation and the modification of ${\bf L}_{N}$
We first observe the impact of different $p$ values on the singular values of matrix $\tilde{\bf F}_{\gamma,N}^{T,p}$. Fig. 1(a) displays the results of the singular values calculated using the SVD functions in Matlab, where $N=200,\gamma=2$ and $T=2$. Similar results are obtained for other combinations of $N$, $\gamma$ and $T$: When $p>0$, some singular values that were originally close to $1$ become decaying, and the decay rate increases with the increase of $p$. When $p$ is too large, the number of singular values greater than the machine precision $\epsilon$ will be significantly reduced. Furthermore, note that the larger $p$ is, the larger the value of $\|\tilde{s}_{\mathcal{L},N}\|$ in the estimate (15) will be, thus affecting the approximation accuracy. In fact, as $p$ increases, we should have chosen a smaller value of $\epsilon$ to achieve approximation, but this is not possible due to the limitation of machine precision.
@@
(b) using ˘$\breve{\mathcal{L}}_{N}$ with $K_{0}=10,C=32$
@@
to conduct a preliminary test on the performance of the algorithm. Here and what follows, the approximation error evaluated as the maximum pointwise error over an equi-sampled grid that is ten times denser than that used for construction. Unless otherwise specified, the value of $\epsilon$ is $141e-$ 14. Figure 2 shows how the error changes with $N$ when the parameters $\omega$, $T$, and $p$ take different values. It is evident that these three parameters significantly affect the calculation accuracy, with their increase leading to a reduction in accuracy. This is because $\exp(\text{i}\pi x)$ corresponds to $\phi_{k}$ when $k=\omega T$, so we can obtain the estimate of $\|\tilde{s}_{\mathcal{L},N}\|$ in (15) as:
$\|\tilde{s}_{\mathcal{L},N}\|\sim\left(\omega T\right)^{p}.$ (23)
@@
˘$\breve{\mathcal{L}}_{N}:\mathbb{C}^{2N+1}\rightarrow\mathbb{C}^{2N+1},{c}_{N}=\{\hat{c}_{k}\}_{|k|\leq N}\mapsto\tilde{c}_{N}$ (24)
with
$\left(\tilde{c}_{N}\right)_{k}=\left\{\begin{aligned} \hat{c}_{k},&&|k|<K_{0},\\ {\left(1+\frac{|k|-K_{0}}{C}\right)^{p}}\hat{c}_{k},&&|k|\geq K_{0},\end{aligned}\right.$ (25)
where $K_{0}$ and $C$ are constants to be determined. Such corrections are reasonable because the goal of introducing the weighted generalized inverse is essentially to characterize the attenuation of the coefficients, which is an asymptotic property when $k$ tends to infinity. The above $K_{0}$ can be estimated during the calculation process, as follows: First, calculate ${\bf c}_{{\bf L},N}^{\epsilon}$ according to (21)–(22) and then select $K_{0}$ as
$K_{0}=\inf\left\{K~{}|\left|\left({\bf c}_{{\bf L},N}^{\epsilon}\right)_{k}\right|>\varepsilon,\forall\,\,|k|>K\right\}.$ (26)

### v3

@@ Abstract
This paper proposes a novel approach for Fourier extension by reformulating the problem as solving a compact operator equation with perturbed data. To mitigate oscillations in the extended region, we introduce a weighted generalized inverse framework that shifts the approximation target from the best approximate solution to a weighted variant. Theoretical error bounds for the solution are established. We identify computational challenges arising from machine precision limitations when implementing the original weighted operator and propose a corrected operator to address this issue. The relevant parameters involved in the method are further tested, and finally the effectiveness of the method is verified through numerical experiments.
@@
Fourier ExtensionsFourier ContinuationWeighted Generalized Inverse
MSC:
42A10 41A17 65T40 42C15
@@ 1 Introduction
Due to its flexibility, frame approximation has received increasing attention in recent years. Its related theories and numerical algorithms are being widely studied and have made substantial progress adcock2019frames ; adcock2020frames ; adcock2020approximating ; adcock2021frame ; herremans2024efficient ; ebner2023regularization . Fourier extension is an important form of frame approximation, which has significantly contributed to the advancement of the field of frame theory. Initially proposed to mitigate the Gibbs phenomenon observed in classical Fourier series, Fourier extension has since witnessed extensive research. In recent years, many scholars have made substantial contributions to developing its mathematical foundations, devising efficient algorithms, reduce artifacts, enhance the resolution of inversion results and exploring its applications across various fields Boyd2002A ; Huybrechs2010 ; Lyon2011 ; Matthysen2016FAST ; 2012Sobolev ; Matthysen2018 ; Bruno2022 ; Webb2020 .
Similar to general frame approximations, Fourier extension also encounters challenges due to its near-redundancy. Specifically, it faces the issue of ill-conditioning in the least-squares system used to compute the coefficients of the approximation. Fortunately, accurate Fourier extension approximation can still obtained via suitable regularization. Currently, Truncated Singular Value Decomposition (TSVD) is the commonly used regularization technique, and various fast algorithms have been developed on this basis Lyon2011 ; Matthysen2016FAST ; zhao2024fast . However, solutions obtained by existing methods exhibit oscillations in the extended region, which is detrimental to the application of the extension method in solving partial differential equations (PDEs). A post-processing method described in the literature 2012Sobolev demonstrates that it is possible to find extension functions with desirable properties without compromising accuracy.
According to the extension theorem of Sobolev spaces, it is possible to identify extension functions with desirable properties. This raises a question: Can we find an effective method to directly calculate such extension functions? In fact, when $N$ is finite, $\{e^{\text{i}\frac{k\pi}{T}t}\}_{k=-N}^{N}$ is a basis on [-1,1], that is, the solution to the least squares problem is unique. However, when the TSVD is used, since the singular vectors with singular values $\sigma_{k}\leq\epsilon$ (where $\epsilon$ is the truncation parameter) are discarded, the problem actually becomes a multi-solution. TSVD only yields the solution which minimizes the Euclidean norm of the Fourier coefficients. When a function has a certain smoothness, its Fourier coefficient has a corresponding decay. The smoother the function, the faster the decay rate. Therefore, it is natural to restrict its decay in the selection of solutions. The idea of enforcing special features of regularization approximations has a long history. In his original paper tikhonov1963regularization , Tikhonov (1963) defined the regularization solution $x_{\alpha}$ of the equation
@@
where $\mathcal{L}$ is a differential operator. Along similar lines, Hansen et. al developed GTSVD and MTSVD methods hansen1989regularization ; hansen1992modified ; hansen1994regularization and Natterer (1984) introduced “regularization in Hilbert scales” natterer1984error . Interestingly, Natterer (1984) has shown that “oversmoothing” (the unknown solution fails to have the expected smoothness) makes no trouble, that is, we can impose stronger constraints than the smoothness of the solution.
@@ 2.1 A new perspective on Fourier extension
From Huybrechs2010 , this system is a tight frame with frame bound $T$. For $f\in H^{s}(\Lambda_{1})$, we consider the compact operator equation
@@
  • From the classical regularization theory, we know that the regularization parameter should be related to the error level $\|f-f_{N}\|_{L^{2}(\Lambda_{1})}$, which explains “the coefficient norm first increases exponentially fast, before decaying to $\mathcal{O}(1)$ in the limit.” mentioned in adcock2021frame . When $N$ is small, the error $\|f-f_{N}\|_{L^{2}(\Lambda_{1})}$ is large, and $\epsilon$ is too small for the regularization parameter required for stable solution. As $N$ increases, the error gradually decreases, and after a certain threshold, $\epsilon$ can meet the requirements of stable solution.Therefore, we can actually draw a conclusion: the solutions with large coefficients are unimportant because they mean that $\|f-f_{N}\|_{L^{2}(\Lambda_{1})}$ is large, that is, a desirable approximation to $f$ has not been obtained.
@@
    $\|f-\mathcal{Q}_{N,1}^{\epsilon}f\|\leq\|f-f_{N}\|_{L^{2}(\Lambda_{1})}+\epsilon\|{s}_{N}\|_{\ell^{2}}$
@@
    $\|f-\mathcal{Q}_{N,2}^{\epsilon}f\|\leq\|f-f_{N}\|_{L^{2}(\Lambda_{1})}+\sqrt{\epsilon}\|{s}_{N}\|_{\ell^{2}}.$
    That is to say, if we solve the normal equation, we can only obtain an accuracy of $O(\sqrt{\epsilon})$ instead of $O({\epsilon})$. The literature adcock2019frames also shows that this is true for general frame approximations. Therefore, we should directly focus on the discretization of equation (2) rather than its normal equation.
@@ 2.2 Fourier extension towards weighted best-approximation solution
Because $f\in H^{s}(\Lambda_{1})$, we can expect the coefficient $\{\hat{c}_{k}\}$ has a corresponding decay property. Therefore, the following weighted best-approximation solution ${c}^{{\dagger}}_{\mathcal{L}_{p}}$ is a more reasonable goal:
@@
$(\mathcal{L}_{p}{c})_{k}=\rho_{k}\hat{c}_{k}$ (9)
with
$\rho_{k}=:\left\{\begin{aligned} &\sqrt{1+|k|^{2p}},&&0\leq p<\infty,\\ &e^{|k|},&&p=\infty.\end{aligned}\right.$
$p=\infty$ corresponds to the case where $f$ is analytic and it can be seen that when $p=0$, ${c}^{{\dagger}}={c}^{{\dagger}}_{\mathcal{L}_{p}}.$
@@
  • Ideally, we would like to take $p=s$, that is, we hope that the solution to the problem has the same smoothness as the function $f$. However, in practice, the value of $s$ may be unknown. In fact, the above operator $\mathcal{L}_{p}$ faces the difficulty of numerical realization due to the limitation of machine precision. We will discuss this in section 3.2 and propose a modified operator.
  • If $p>s$, that is , the smoothness of the solution is lower than the requirement of the operator $\mathcal{L}_{p}$, i. e., no solution of (1) is an element of $\mathcal{D}(\mathcal{L}_{p})$, this is called oversmoothing. The results in natterer1984error ; yu2024super have shown that the regularization method can still obtain the optimal convergence order in the case of oversmoothing (when the regularization parameters are reasonably selected).
Now we set the approximation target the weighted generalized solution ${c}^{{\dagger}}_{\mathcal{L}}$. At this time, the approximation of the solution can be achieved by the “regularization in the Hilbert scale”natterer1984error . In fact, if we let
$\psi_{k}(x)=\frac{1}{\rho_{k}}\phi_{k}(x),\quad\hbox{and}\quad\Psi=\{\psi_{k}\}_{k\in\mathbb{Z}},$
@@
${c}^{{\dagger}}_{\mathcal{L}_{p}}=\mathcal{L}_{p}^{-1}\tilde{c}^{{\dagger}}.$
@@
$\mathcal{L}_{p,N}:\mathbb{C}^{2N+1}\rightarrow\mathbb{C}^{2N+1},{c}_{N}=\{\hat{c}_{k}\}_{|k|\leq N}\mapsto\tilde{c}_{N}=\{\rho_{k}\hat{c}_{k}\}_{|k|\leq N},$
@@
~$\tilde{\mathcal{F}}_{N}=\mathcal{F}_{N}\mathcal{L}^{-1}_{p,N},\quad{s}_{\mathcal{L},N}=\mathcal{P}_{N}{c}_{\mathcal{L}_{p}}^{{\dagger}}\quad\tilde{s}_{\mathcal{L},N}=\mathcal{L}_{p,N}{s}_{\mathcal{L},N},\quad f_{\mathcal{L},N}=\tilde{\mathcal{F}}_{N}\tilde{s}_{\mathcal{L},N}.$
@@
$c_{\mathcal{L},N}^{\epsilon}=\mathcal{L}_{p,N}^{-1}\tilde{c}_{N}^{\epsilon}$ (13)
@@
$\|f-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f\|_{L^{2}(\Lambda_{1})}\leq\|f-f_{\mathcal{L,N}}\|_{L^{2}(\Lambda_{1})}+\epsilon\|\tilde{s}_{\mathcal{L},N}\|_{\ell^{2}}.$ (15)
Furthermore, if $f\in H^{s}(\Lambda_{1})$, then we have
$\|f-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f\|_{L^{2}(\Lambda_{1})}\leq C\left(N^{-s}+\epsilon\right)\|f\|_{H^{s}(\Lambda_{1})},$ (16)
where $C>0$ is independent of f and N.
Proof
@@
From the the extension theorem of Sobolev spaces, we know that
$\|\tilde{s}_{\mathcal{L},N}\|_{\ell^{2}}\leq C\|f\|_{H^{s}(\Lambda_{1})}.$
In particular,
$\|f-f_{\mathcal{L,N}}\|_{L^{2}(\Lambda_{1})}\leq\|f-f_{\mathcal{L,N}}\|_{L^{2}(\Lambda_{2})}\leq CN^{-s}\|f\|_{H^{s}(\Lambda_{1})}$
Therefore,
$\|f-{\mathcal{Q}}_{\mathcal{L},N}^{\epsilon}f\|_{L^{2}(\Lambda_{1})}\leq C\left(N^{-s}+\epsilon\right)\|f\|_{H^{s}(\Lambda_{1})}.$
@@ 3.1 Discretization based on uniform sampling
Selecting different discrete nodes will affect the discrete accuracy of the operator $\mathcal{F}_{N}$, thus having a substantial impact on the calculation results. Previous work distinguishes between uniform sampling, and an optimal sampling set called mapped symmetric Chebychev nodes adcock2014resolution . In this paper, we focus on the case where the sampling points are uniform:
@@
$\left({\bf F}_{\gamma,N}^{T}\right)_{\ell,k}=\frac{1}{\sqrt{L}}\phi_{k}(t_{\ell})$
and ${\bf L}_{p,N}$ be a diagonal matrix with
$({\bf d})_{k}={\rho_{k}},\quad k=-N,\ldots,N.$ (19)
@@
$\tilde{\bf F}_{\gamma,N}^{T,p}={\bf F}_{\gamma,N}^{T}{\bf L}_{p,N}^{-1}$ (20)
@@
${\bf c}_{{\bf L},N}^{\epsilon}={\bf L}_{p,N}^{-1}\tilde{\bf c}_{N}^{\epsilon},\quad{\bf Q}_{{\bf L},N}^{\epsilon}{\bf f}={\bf F}_{\gamma,N}^{T,p}{\bf c}_{{\bf L},N}^{\epsilon}.$ (23)
@@ 3.2 The influence of the parameter $p$ on calculation and the modification of ${\bf L}_{N}$
We first observe the impact of different $p$ values on the singular values of matrix $\tilde{\bf F}_{\gamma,N}^{T,p}$. Fig. 1(a) illustrates the results of the singular values calculated using the SVD functions in Matlab, where $N=200,\gamma=2$ and $T=2$. Similar results are obtained for other combinations of $N$, $\gamma$ and $T$: When $p>0$, some singular values that were originally close to $1$ become decaying, and the decay rate increases with the increase of $p$. When $p$ is too large, the number of singular values greater than the machine precision $\epsilon$ will be significantly reduced. Furthermore, note that the larger $p$ is, the larger the value of $\|\tilde{s}_{\mathcal{L},N}\|$ in the estimate (15) will be, thus affecting the approximation accuracy. In fact, as $p$ increases, we should have chosen a smaller value of $\epsilon$ to achieve approximation, but this is not possible due to the limitation of machine precision.
@@
(b) using ˘$\breve{\mathcal{L}}_{N}$ with $K_{0}=20$
@@
to conduct a preliminary test on the performance of the algorithm. Here and what follows, the approximation error evaluated as the maximum pointwise error over an equi-sampled grid that is ten times denser than that used for construction. Unless otherwise specified, the value of $\epsilon$ is $141e-$ 14. Fig. 2 shows how the error changes with $N$ when the parameters $\omega$, $T$, and $p$ take different values. It is evident that these three parameters significantly affect the calculation accuracy, with their increase leading to a reduction in accuracy. This is because $\exp(\text{i}\pi\omega x)$ corresponds to $\phi_{k}$ when $k=\omega T$, so we can obtain the estimate of $\|\tilde{s}_{\mathcal{L},N}\|$ in (15) as:
$\|\tilde{s}_{\mathcal{L},N}\|_{\ell^{2}}\sim\left\{\begin{aligned} &\left(\omega T\right)^{p},&0\leq p<\infty,\\ &e^{\omega T},&p=\infty.\end{aligned}\right.$ (24)
@@
˘$\breve{\mathcal{L}}_{N}:\mathbb{C}^{2N+1}\rightarrow\mathbb{C}^{2N+1},{c}_{N}=\{\hat{c}_{k}\}_{|k|\leq N}\mapsto\tilde{c}_{N},$ (25)
when $0\leq p<\infty$
$\left(\tilde{c}_{N}\right)_{k}=\left\{\begin{aligned} &\frac{1}{K_{0}}\hat{c}_{k},&&|k|<K_{0},\\ &\left[{\left(\frac{|k|-K_{0}}{(N-K_{0})^{\frac{K_{0}}{K_{0}+C}}T}\right)^{p}}+\frac{1}{K_{0}}\right]\hat{c}_{k},&&|k|\geq K_{0},\end{aligned}\right.$ (26)
and when $p=\infty$
$\left(\tilde{c}_{N}\right)_{k}=\left\{\begin{aligned} &\frac{1}{K_{0}}\hat{c}_{k},&&|k|<K_{0},\\ &\left[\exp\left(\frac{|k|-K_{0}}{(N-K_{0})^{\frac{K_{0}}{K_{0}+C}}T}\right)-\frac{K_{0}}{K_{0}+1}\right]\hat{c}_{k},&&|k|\geq K_{0},\end{aligned}\right.$ (27)
where $K_{0}$ and $C$ are constants to be determined. Such corrections are reasonable because the goal of introducing the weighted generalized inverse is essentially to characterize the attenuation of the coefficients, which is an asymptotic property when $k$ tends to infinity. The parameter $K_{0}$ describes the starting point of the function attenuation, which can be estimated during the calculation process, as follows: First, replace ${\bf L}_{N}$ with $\breve{{\bf L}}_{N}(p=4,K_{0}=0)$ in (22)–(23) to calculate ${\bf c}_{\breve{\bf{L}},N}^{\epsilon}$ and then select $K_{0}$ as
$K_{0}=\inf\left\{K~{}|\left|\left({\bf c}_{\breve{\bf L},N}^{\epsilon}\right)_{k}\right|>\varepsilon,\forall\,\,|k|>K\right\}.$ (28)